OpenStudy (anonymous):

The number of people in a town of 10,000 who have heard a rumor started by a small group of people is given by the following function: N(t)=10,000/5+1245e^-.97t How long will it be until 1,000 of the people in the town have heard the rumor?

OpenStudy (mertsj):

Replace N(t) with 1000 and solve for t

OpenStudy (anonymous):

@Mertsj thanks, i got t=-2.1188

OpenStudy (anonymous):

then if i plug that back in i get 11721.8072 is that the final answer?

OpenStudy (mertsj):

Your answer should not be negative. The question asks how long it will be. So t which is time, is the answer to the question the problem asks.

OpenStudy (anonymous):

can you please tell me what it is?

OpenStudy (mertsj):

Why does the problem have 10000/5 instead of 2000?

OpenStudy (mertsj):

\[N(t)=\frac{10,000}{5}+1245e ^{(-.97t)}\]

OpenStudy (mertsj):

Is that the given equation?

OpenStudy (anonymous):

yes that is correct @Mertsj

OpenStudy (anonymous):

no sorry that is wrong

OpenStudy (anonymous):

the equation is:

OpenStudy (mertsj):

\[5000+1245000e ^{-.97t}=10000\] \[1245000e ^{-.97t}=5000\] \[e ^{-.97t}=.0040\] \[-.97t \ln e=\ln .0040\] \[-.97t=-5.517\] t=5.688

OpenStudy (anonymous):

thank you so much, so in that case how do i find (How long will it be until 1,000 of the people in the town have heard the rumor?)

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

What does t represent?

OpenStudy (anonymous):

i dont know :(

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Ah ha. That is your problem.

OpenStudy (mertsj):

Did you post the entire problem or was there more words that you didn't include?

OpenStudy (mertsj):

t is the time. When the problem says "how long will it be" it is asking you for the time. So now we have found that it will be 5.688 hours? days? until 1000 people have heard the rumor. The time units are not given unless there is more to the problem that you didn't post.

OpenStudy (anonymous):

that is the entire problem, it must be 5 hours then because there is no other information given @Mertsj